OpenStudy (anonymous):

Plz help i will fan Which two temperatures have a sum of 0 degrees Celsius? 2 degrees Celsius and -2 degrees Celsius 2 degrees Celsius and 2 degrees Celsius -2 degrees Celsius and -2 degrees Celsius 2 degrees Celsius and 0 degrees Celsius

OpenStudy (daniellelovee):

a negative and a positive of the same value will always equal 0

OpenStudy (anonymous):

.what is 2-2=? 2+2=? -2+2=? 2+0=?

OpenStudy (daniellelovee):

no it says there is a sum @magepker728 so it should be 2+-2=? 2+2=? -2+-2=? 2+0=

OpenStudy (anonymous):

so it's the first one! @Daniellelovee

OpenStudy (daniellelovee):

I think so :)

OpenStudy (anonymous):

Ok thanks for the help! @Daniellelovee I'll fan you

OpenStudy (daniellelovee):

ok thank you :)

OpenStudy (anonymous):

and 3rd one

OpenStudy (daniellelovee):

no because -2+-2=-4